A 9.00 m × 14.0 m swimming pool slopes linearly from a 1.10 m depth at one end to a 3.00 m depth at the other.

Part A

What is the mass of water in the pool?

Express your answer with the appropriate units.

=kg
Answer #1

Volume of the pool = 9*14*1.1 + 0.5*14*(3 - 1.1)*9

= 258.3 m^3

Mass of water = volume*density

= 258.3*1000

= 2.583*10^5 kg <<<<<<------Answer